OpenStudy (megatron):

Find the solutions of g(x). Show each step. g(x) = x^2 +6x + 1 If you could show me how to do this so I can do it in the future, it would be greatly appreciated. (:

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} g(x) = x^2 &+6x &+ 1\\ &\uparrow &\uparrow \\ &3-2&3\cdot -2 \end{array}\)

OpenStudy (jdoe0001):

woops, darn the went wrong =) one sec

OpenStudy (jdoe0001):

I mistook the equation... well I gather you'd need to use the quadratic formula for this one

OpenStudy (jdoe0001):

since you can't squeeze 6 out of 1

OpenStudy (jdoe0001):

unless you mean \(\bf x^2+1x+6\)

OpenStudy (jdoe0001):

well, not even that one... would work well... unless you meant say \(\bf x^2-1x-6\)

OpenStudy (megatron):

No the equation is g(x)=x^2+6x+1

OpenStudy (jdoe0001):

I'm assuming you're meant to use the perfect square trinomial to solve for "x"?

OpenStudy (megatron):

My instructions just say to find the solutions of g(x). The title of the assignment is completing the square, if that helps..

OpenStudy (jdoe0001):

anyhow...lemme do that instead.... lemme make it a perfect square trinomial anyhow... to solve quadratics for "x", you'd set y=0, that is g(x) = x^2 +6x + 1 => 0= x^2 +6x + 1 thus \(\bf g(x) = x^2 +6x + 1\implies g(x) = (x^2 +6x+{\color{red}{ 3}}^2) + 1-{\color{red}{ 3}}^2 \\ \quad \\ g(x)=(x+3)^2+1-9\implies g(x)=(x+3)^2-8\quad \textit{setting y=0} \\ \quad \\ 0=(x+3)^2-8\implies 8=(x+3)^2\qquad taking\quad \sqrt{\qquad } \\ \quad \\ \pm\sqrt{8}=x+3\implies 2\pm\sqrt{2}=x+3\implies 2\pm\sqrt{2}-3=x\)

OpenStudy (megatron):

So the final answer would be 2 plus or minus radical 2 minus 3?

OpenStudy (jdoe0001):

after completing the square and solving for "x", yes is a QUADRATIC, thus it has 2 answers